OpenStudy (anonymous):

could you help I am doing some home work and i cant get these 2 8. y = –x + 2 y = 3x – 1 13. –4x + 3y = –12 –2x + 3y = –18

OpenStudy (nottim):

substitution?

OpenStudy (anonymous):

for the first one, set them equal and solve

OpenStudy (ganpat):

y = –x + 2 y = 3x – 1 yes, moreover equate both y... so, 3x-1 = -x +2 3x+x = 2 +1 4x = 3 x =3/4... and y = -x+2 = -3/4 + 2 = 5/4...

OpenStudy (ganpat):

For the second one, subtract equation 1 & 2.. such that y gets cancelled and solve for x.. after you find x, substitute in any of the equation and find y...

OpenStudy (anonymous):

thx

OpenStudy (ganpat):

W/c .. :)